HUBBLE SPACE TELESCOPE DAILY REPORT #2976 PERIOD COVERED: 0000Z (UTC) 10/12/01 - 0000Z (UTC) 10/15/01 Daily Status Report as of 288/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Six Sets of STIS/CCD 8901 (Dark Monitor-Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. There was no anomalous activity. 1.2 Completed Seven Sets of WF/PC-2 8936 (Cycle 10 Supplemental Darks Pt1/3) The WF/PC-2 was used to perform a dark calibration program that obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. As described in HSTAR 8367 and 2.1, the acquisition for the sixth and seventh iterations of this proposal failed, the take data flag remained down, and observations were lost. Otherwise, the proposal completed with no other reported problems. 1.3 Completed STIS/CCD/MA2 8673 (The Properties of Ly-Alpha Absorbers at Redshifts Between 0.9HST) The Space Telescope Imaging Spectrograph (CCD and MA2) was used obtain new Echelle resolution spectra of three bright quasars which, when combined with archived and scheduled observations, will be used to characterize the properties of Ly-Alpha absorbers in the redshift range 0.9